# Generative AI for Cybersecurity of Energy Management Systems: Methods, Challenges, and Future Directions

### v1

@@ 1.5 Paper Structure
The rest of this paper is organized as follows: Section 2 presents a multi-point attack model considering the EMS workflow, before and after the SE process. Section 3 demonstrates the comprehensive description of different attack points, their modeling, miscellaneous scenarios, and the GPT implementation with some direct responses using the trained GPT model within the power system domain. Finally, conclusions and directions for future work are outlined in Section 4.
@@
2 A Proposed Multi-point Attack Model in EMSs
@@
Different attack points based on stealth attacks (i.e., attack point #1) and intentional/unintentional attacks (i.e., attack points #2 and #3) including cyberattacks, system errors, and FDI attacks are represented in Fig. 1.
@@
The modern EMS faces sophisticated cyber threats that exploit vulnerabilities at different stages of the data processing pipeline. Understanding these attack vectors is crucial for developing comprehensive defense mechanisms that protect the integrity of power system operations [39]. To recap, Table 2 demonstrates these attack vectors with their proposed AD solutions.
Table 2: Different attack points and their proposed ADSs.
@@
1 Stealth attacks GenAI-based AD
2 FDI attacks GenAI-based AD
3 HMI screen corruption attacks/errors SoM-GI-based AD
@@
3 Attack Vectors and Experimental Validation
3.1 Attack Point #1: Stealth Attacks (Gray Dashed Line)
@@ A Mathematical Construction of SE and Stealth Attacks
where $\mathbf{H}_{k}^{\dagger}=(\mathbf{H}_{k}^{T}R^{-1}\mathbf{H}_{k})^{-1}\mathbf{H}_{k}^{T}R^{-1}$ and $\mathbf{H}_{k}$ represent the Jacobian matrix of $\mathbf{H}$ evaluated at iteration $k$. Under Gaussian noise assumptions, this WLS approach yields optimal results. Upon convergence, achieved when $\|\hat{\mathbf{x}}_{k}-\hat{\mathbf{x}}_{k-1}\|<\delta$ for a predefined small threshold $\delta>0$—the resulting residuals experience analysis to identify potential measurement anomalies through verification of Gaussian properties. Such anomalies in data might stem from natural failures (e.g., sensor malfunctions or communication disruptions, or potentially from deliberate FDI attacks). Standard detection methods typically employ $\chi^{2}$ testing for identifying anomalous data. Additionally, a DC SE variant exists where only phase angles require estimation, with voltages assumed to be at unity (1 p.u.). This simplified model neglects line resistances and assumes small phase angle differences between buses, resulting in a linear regression framework as Eq. (3[22]:
@@
This characteristic of the mathematical model enables the attack approach to evade conventional mechanisms designed for detecting erroneous data, as $J(\mathbf{x}_{a})=J(\mathbf{x})\leq\tau$ [20]. According to a case study for an IEEE 14-bus system, 300 attack points are applied to the system at each bus to find the range of stealth attacks for different buses, forced by the NERC regulation in terms of bus voltage magnitude violation, which states that the bus voltages should fall within the range of 0.95–1.05 p.u. to meet the requirements [42]. According to this process, a sample of the stealth attack range for Bus 2 is illustrated in Table 3. All these attack points are also injected to all other buses to find the range of stealth attack as illustrated in Fig. 2.
Table 3: A part of attack points applied to Bus 2 to find the stealth attack range.
@@
The presented visualization reveals the feasible ranges across the IEEE 14-bus system that successfully evade the $\chi^{2}$ as a BDD system while maintaining system observability. Further, a numerical representation of stealth attack ranges including the start point, end point, width of the range, and original bus voltages based on the numerical representation is given in Table 4.
Table 4: Stealth attack ranges results for an IEEE 14-bus system based on start/end points of attacks along with the original voltage before applying the NERC regulation.
@@
3.1.1 Scenario #1A: 5-Point Distributed Stealth Attack
This scenario represents a 5-point distributed stealth attack designed to evade the $\chi^{2}$ test based on the BDD while maintaining all PF constraints. The attack strategically targets five measurement points across the IEEE 14-bus system with coordinated changes as given in the attack vector (11).
@@ 3.1.1 Scenario #1A: 5-Point Distributed Stealth Attack
3.1.2 Scenario #1B: 8-Point Massive Coordinated Attack
This scenario demonstrates an even more complex 8-point massive coordinated attack that exploits the fundamental limitations of statistical BDD through measurement manipulation. This attack simultaneously modifies eight critical measurements across the IEEE 14-bus system as represented in the attack vector (12).
@@ 3.1.2 Scenario #1B: 8-Point Massive Coordinated Attack
Additionally, power measurements at buses $\{2,4,5,10,12,13,14\}$ show small decreases of 0.0357 p.u. each.
State Estimation Analysis
Based on the Power Flow Solution:
@@
    $P_{i}$ $=V_{i}\sum_{j}V_{j}(G_{ij}\cos(\theta_{i}-\theta_{j})+B_{ij}\sin(\theta_{i}-\theta_{j}))\quad\checkmark$
    $Q_{i}$ $=V_{i}\sum_{j}V_{j}(G_{ij}\sin(\theta_{i}-\theta_{j})-B_{ij}\cos(\theta_{i}-\theta_{j}))\quad\checkmark$
@@
GenAI Analysis and Detection
The correlation analysis reveals:
@@
State Estimation Analysis
based on the Power Flow Solution:
@@
GenAI Analysis and Detection
GenAI’s comprehensive analysis reveals severe anomalies:
@@
$\text{Correlation }V_{2}\text{-}P_{2}:\quad w$ $=0.73\text{ (abnormal)}$
$\text{Correlation }P_{4}\text{-}P_{13}:\quad w$ $=0.81\text{ (coordinated opposites)}$
$\text{Correlation }V_{4}\text{-}V_{11}:\quad w$ $=0.68\text{ (opposing pattern)}$
@@
Chi-Square Test Perspective
For both scenarios, chi-square concludes NORMAL OPERATION because:
@@
Semantic Understanding
GenAI identifies both scenarios as attacks through:
@@
3.2 Attack Point 2: Post-State Estimation Attack in EMS Database (Red Dashed Line)
@@
3.2.1 Scenario #2A: State Vector Manipulation Attack
@@
$\text{Post-SE Output: }\hat{\mathbf{x}}=[V_{1},\theta_{1},P_{1},Q_{1},V_{2},\theta_{2},P_{2},Q_{2},\ldots,V_{n},\theta_{n},P_{n},Q_{n}]^{T}$ (13)
@@
$\hat{\mathbf{x}}_{\text{corrupted}}=\hat{\mathbf{x}}+\Delta\mathbf{x}\quad\text{Where}\quad\Delta\mathbf{x}=[\Delta{V_{1}},\Delta\theta_{1},\Delta{P_{1}},\Delta{Q_{1}},\Delta{V_{2}}\ldots]^{T}$ (15)
$\text{Sample Result in EMS Database: }P^{\prime}_{ij}=|V_{i}+\Delta{V_{i}}||V_{j}+\Delta{V_{j}}||Y_{ij}|\sin(\theta_{i}-\theta_{j}+\Delta\theta_{i}-\Delta\theta_{j})$ (16)
@@ 3. System-Wide Impact
3.2.2 Scenario #2B: Topology Corruption Attack
@@ 3. Stress Indicators
3.2.3 Complex Scenarios for Attack Point 2
In order to show the better capability of the GenAI-based AD in detecting and distinguishing normal and abnormal scenarios, a contingency analysis was performed in PowerWorld Simulator for an IEEE 14-Bus system to generate normal scenarios. These generated normal scenarios show the different statuses of normal PF datasets that help the GenAI to learn about the contingency analysis as well as training the GenAI with SE and PF rules. Hence, a series of 30 scenarios according to the contingency analysis were generated as shown in Table 5. These scenarios are defined based on a combination of CB statuses, transformer tap changes, load changes, and power limits to include different types of cases.
Table 5: Power system test scenarios implemented in PowerWorld Simulator for the GenAI tool training.
@@ 3.2.3 Complex Scenarios for Attack Point 2
The key learning here is that state estimation can identify extreme but valid operating conditions. The progression from normal operation $\rightarrow$ partial islanding $\rightarrow$ complete islanding demonstrates how the system can evolve under severe contingencies while maintaining physical validity. This is a valuable example of why understanding state estimation theory and physical principles is crucial for proper anomaly detection, rather than simply pattern matching against known scenarios. As can be seen, the open CB between Bus 9 and Bus 10 is correctly detected as a normal condition in an islanding mode with other relevant information according to the data pattern and SE rules.
@@
3.3 Attack Point 3: Attack/Error in HMI Screen based on RTDB (Purple Dashed Line)
@@
3.3.1 A Normal HMI Screen with Power Flow Analysis Using Only GenAI Without SoM Technique
The following “Prompt”-“Response” boxes show a normal HMI screen without any attack/error in segmented parts. Then, a prompt requested the GPT to arrange the segments based on its understanding. The process is that the 9 segments are given as inputs along with the following prompt, without any additional rules or principles to train the GPT model. As can be seen, the response demonstrates the incapability of the GPT model to arrange and make a connection between different segments of an HMI screen based on the 14-bus system in PowerWorld Simulator. {promptbox} Scenario #3A (GenAI without SoM)

### v2

@@ I-F Paper Structure
The rest of this paper is organized as follows: Section II presents a multi-point attack model considering the EMS workflow, before and after the SE process. Section III demonstrates the comprehensive description of different attack points, their modeling, miscellaneous scenarios, and the GPT implementation with some direct responses using the trained GPT model within the power system domain. Finally, conclusions and directions for future work are outlined in Section IV.
@@
II A Proposed Multi-point Attack Model in Energy Management Systems
@@
Different attack points based on stealth attacks (i.e., attack point #1) and intentional/unintentional attacks (i.e., attack points #2 and #3) including cyberattacks, system errors, and FDI attacks, are represented in Fig. 1.
@@
Prior to delving into the detailed analysis of individual attack vectors, it is essential to clarify that GenAI assumes a comprehensive, integrated role within the architecture of the framework. This functionality stands in contrast to the simplistic notion of GenAI operating as a tool with a singular purpose. First, it operates as a language model that interprets natural language descriptions of system states, operational rules, and component relationships, enabling operators and engineers to interact with the detection system using familiar terminology rather than requiring specialized query languages. Second, it functions as a multimodal reasoning engine that synthesizes information from diverse sources including numerical measurements from state estimators, visual representations in HMI displays, and textual operational logs, to identify inconsistencies that span multiple data modalities. Third, it acts as an anomaly classifier that evaluates whether observed conditions violate physical laws, operational constraints, or expected behavioral patterns, subsequently categorizing detected anomalies into specific threat classes such as stealth attacks, database manipulations, or display corruptions. These roles are not implemented as separate modules but rather emerge inherently from GenAI’s capabilities when properly guided through structured prompts and domain-specific knowledge. This integrated approach distinguishes the suggested framework from conventional AD systems that typically address either numerical analysis or visual inspection, but rarely within a unified reasoning process.
@@
The modern EMS faces sophisticated cyber threats that exploit vulnerabilities at different stages of the data processing pipeline. Understanding these attack vectors is crucial for developing comprehensive defense mechanisms that protect the integrity of power system operations [47]. To recap, Table III demonstrates these attack vectors with their proposed AD solutions.
TABLE III: Different attack points and their proposed ADSs.
@@
1
Stealth attacks
GenAI-based AD
2
FDI attacks
GenAI-based AD
3
HMI screen corruption
attacks/errors
SoM-GI-based AD
@@
II-A GenAI Implementation Methodology and Prompt Engineering
This subsection provides a detailed explanation of how GenAI technology is implemented within the AD framework, addressing the fundamental question of whether LLMs are fine-tuned on power system data or employed through structured prompting methodologies.
II-A1 Model Selection and Deployment Strategy
This proposed framework utilizes Anthropic Claude Pro [39], a state-of-the-art LLM, in its pre-trained configuration without domain-specific fine-tuning. This design choice reflects several strategic considerations that distinguish the approach from traditional ML methodologies commonly applied to power system security. Rather than fine-tuning the model on extensive EMS-specific datasets, an approach that would require collecting, labeling, and curating thousands of operational scenarios and attack patterns, the model’s existing knowledge leverages a base that already encompasses fundamental physics principles, electrical engineering concepts, mathematical reasoning capabilities, and visual semantic understanding. This pre-trained knowledge provides a robust foundation for understanding power system behavior without the time-intensive and resource-demanding process of model fine-tuning. The decision to avoid fine-tuning offers several practical advantages for EMS applications. First, it significantly reduces the deployment timeline from months to weeks, as organizations do not need to collect extensive training datasets before implementation. Second, it maintains the model’s general reasoning capabilities and broad knowledge base, preventing the overfitting issues that can occur when models are narrowly trained on specific datasets. Third, it enables rapid updates to operational rules and constraints by simply modifying prompts rather than retraining models, a critical advantage for adapting to evolving grid configurations and emerging threat landscapes.
II-A2 Structured Prompt Engineering Framework
The core of this methodology lies in systematic prompt engineering that transforms the general-purpose language model into a specialized EMS AD system. These prompts are carefully structured to provide the model with four complementary types of information that collectively enable sophisticated reasoning about power system states and potential anomalies including numerical data integration (i.e., SE, power flow results), power system rules and constraints (e.g., Kirchhoff’s Current Law (KCL)), component descriptions for HMI analysis (e.g., red square markers indicate closed CBs at transmission line terminals), and reference scenarios and in-context learning (ICL) which can be generated using the contingency analysis based on the power flow processes in PowerWorld Simulator. These reference cases serve as benchmark scenarios that guide the model’s interpretation of subsequent test scenarios, effectively creating a learned context without modifying the underlying model parameters. Hence, the prompt-based methodology offers several distinct advantages compared to fine-tuning LLMs on power system-specific datasets, including rapid deployment, transparency and interpretability, easy adaptation, preservation of general knowledge, and reduced data requirements.
To clarify the implementation methodology, this study employs Anthropic Claude Pro as the primary GenAI tool, as referenced in [39]. The simulation process consists of several key steps. Initially, the Claude Pro model was trained using PF results, SE rules, and normal operational scenarios from the IEEE 14-bus system. Then, Excel files containing both normal and attacked data scenarios are given as inputs. Further, domain-specific prompts incorporating power system knowledge and detection rules were developed. The analysis was executed where the trained model examines uploaded data to detect anomalies based on learned patterns and physical constraints. Finally, the framework validated against multiple attack scenarios to verify detection performance. Throughout this paper, the “Implementation in GenAI Tool” subsections serve multiple critical purposes. They provide transparency by demonstrating actual interactions between researchers and the AI system with exact prompts and responses and enable reproducibility by offering sufficient methodological detail, then validate the approach by showing real GenAI outputs rather than theoretical capabilities. Hence, it can bridge the gap between theoretical framework and practical implementation, making the methodology accessible to other researchers in this domain.
III Attack Vectors and Experimental Validation
III-A Attack Point #1: Stealth Attacks (Gray Dashed Line)
@@ A Mathematical Construction of SE and Stealth Attacks
where $\mathbf{H}_{k}^{\dagger}=(\mathbf{H}_{k}^{T}R^{-1}\mathbf{H}_{k})^{-1}\mathbf{H}_{k}^{T}R^{-1}$ and $\mathbf{H}_{k}$ represent the Jacobian matrix of $\mathbf{H}$ evaluated at iteration $k$. Under Gaussian noise assumptions, this WLS approach yields optimal results. Upon convergence, achieved when $\|\hat{\mathbf{x}}_{k}-\hat{\mathbf{x}}_{k-1}\|<\delta$ for a predefined small threshold $\delta>0$, the analysis of resulting residuals detects potential measurement anomalies by verifying Gaussian properties. Such anomalies in data might stem from natural failures (e.g., sensor malfunctions or communication disruptions, or potentially from deliberate FDI attacks). Standard detection methods typically employ $\chi^{2}$ testing for identifying anomalous data. Additionally, a DC SE variant exists where only phase angles require estimation, with voltages assumed to be at unity (1 p.u.). This simplified model neglects line resistances and assumes small phase angle differences between buses, resulting in a linear regression framework as Eq. (3[22]:
@@
This characteristic of the mathematical model enables the attack approach to evade conventional mechanisms designed for detecting erroneous data, as $J(\mathbf{x}_{a})=J(\mathbf{x})\leq\tau$ [20]. According to a case study for an IEEE 14-bus system, 300 attack points are applied to the system at each bus to find the range of stealth attacks for different buses, forced by the NERC regulation in terms of bus voltage magnitude violation, which states that the bus voltages should fall within the range of 0.95–1.05 p.u. to meet the requirements [50]. According to this process, a sample of the stealth attack range for Bus 2 is illustrated in Table IV. All these attack points are also injected to all other buses to find the range of stealth attack as illustrated in Fig. 2.
TABLE IV: A part of attack points applied to Bus 2 to find the stealth attack range.
@@
The presented visualization reveals the feasible ranges across the IEEE 14-bus system that successfully evade the $\chi^{2}$ as a BDD system while maintaining system observability. Further, a numerical representation of stealth attack ranges including the start point, end point, width of the range, and original bus voltages based on the numerical representation is given in Table V.
TABLE V: Stealth attack ranges results for an IEEE 14-bus system based on start/end points of attacks along with the original voltage before applying the NERC regulation.
@@
III-A1 Scenario #1A: 5-Point Distributed Stealth Attack
This scenario represents a 5-point distributed stealth attack (i.e. Eq. (11)) designed to evade the $\chi^{2}$ test based on the BDD while maintaining all PF constraints. The attack strategically targets five measurement points across the IEEE 14-bus system with coordinated changes as given in the attack vector.
@@ III-A1 Scenario #1A: 5-Point Distributed Stealth Attack
III-A2 Scenario #1B: 8-Point Massive Coordinated Attack
This scenario demonstrates an even more complex 8-point massive coordinated attack (i.e. Eq. (12)) that exploits the fundamental limitations of statistical BDD through measurement manipulation.
@@ III-A2 Scenario #1B: 8-Point Massive Coordinated Attack
This attack simultaneously modifies eight critical measurements across the IEEE 14-bus system as represented in the attack vector.
@@
Additionally, power measurements at buses 2,4,5,10,12,13,142,4,5,10,12,\\ 13,14 show small decreases of 0.0357 p.u. each.
State Estimation Analysis Based on the Power Flow Solution:
@@
    $P_{i}$ $=V_{i}\sum_{j}V_{j}(G_{ij}\cos(\theta_{i}-\theta_{j})+B_{ij}\sin(\theta_{i}-\theta_{j}))\qquad\checkmark$
    $Q_{i}$ $=V_{i}\sum_{j}V_{j}(G_{ij}\sin(\theta_{i}-\theta_{j})-B_{ij}\cos(\theta_{i}-\theta_{j}))\qquad\checkmark$
@@
GenAI Analysis and Detection The correlation analysis reveals:
@@
State Estimation Analysis based on the Power Flow Solution:
@@
GenAI Analysis and Detection GenAI’s comprehensive analysis reveals severe anomalies:
@@
$\text{Correlation }V_{2}\text{-}P_{2}:\qquad w$ $=0.73\text{ (abnormal)}$
$\text{Correlation }P_{4}\text{-}P_{13}:\qquad w$ $=0.81\text{ (coordinated opposites)}$
$\text{Correlation }V_{4}\text{-}V_{11}:\qquad w$ $=0.68\text{ (opposing pattern)}$
@@
Chi-Square Test Perspective For both scenarios, chi-square concludes NORMAL OPERATION because:
@@
Semantic Understanding GenAI identifies both scenarios as attacks through:
@@
III-B Attack Point 2: Post-State Estimation Attack in EMS Database (Red Dashed Line)
@@
III-B1 Scenario #2A: State Vector Manipulation Attack
@@
$\text{Post-SE Output: }\hat{\mathbf{x}}=[V_{1},\theta_{1},P_{1},Q_{1},V_{2},\theta_{2},P_{2},Q_{2},\ldots,V_{n},\theta_{n},P_{n},Q_{n}]^{T}$
@@
$\hat{\mathbf{x}}_{\text{corrupted}}=\hat{\mathbf{x}}+\Delta\mathbf{x}\quad\text{Where}\quad\Delta\mathbf{x}=[\Delta{V_{1}},\Delta\theta_{1},\Delta{P_{1}},\Delta{Q_{1}},\Delta{V_{2}}\ldots]^{T}$
$P^{\prime}_{ij}=|V_{i}+\Delta{V_{i}}||V_{j}+\Delta{V_{j}}||Y_{ij}|\sin(\theta_{i}-\theta_{j}+\Delta\theta_{i}-\Delta\theta_{j})$
@@ 3. System-Wide Impact
III-B2 Scenario #2B: Topology Corruption Attack
@@ 3. Stress Indicators
III-B3 Complex Scenarios for Attack Point 2
In order to show the better capability of the GenAI-based AD in detecting and distinguishing normal and abnormal scenarios, a contingency analysis was performed in PowerWorld Simulator for an IEEE 14-Bus system to generate normal scenarios. These generated normal scenarios show the different statuses of normal PF datasets that help the GenAI to learn about the contingency analysis as well as training the GenAI with SE and PF rules. Hence, a series of 30 scenarios according to the contingency analysis were generated as shown in Table VI. These scenarios are defined based on a combination of CB statuses, transformer tap changes, load changes, and power limits to include different types of cases.
TABLE VI: Power system test scenarios implemented in PowerWorld Simulator for the GenAI tool training.
@@ III-B3 Complex Scenarios for Attack Point 2
The key learning here is that state estimation can identify extreme but valid operating conditions. The progression from normal operation $\rightarrow$ partial islanding $\rightarrow$ complete islanding demonstrates how the system can evolve under severe contingencies while maintaining physical validity. This is a valuable example of why understanding state estimation theory and physical principles is crucial for proper AD, rather than simply pattern matching against known scenarios. As can be seen, the open CB between Bus 9 and Bus 10 is correctly detected as a normal condition in an islanding mode with other relevant information according to the data pattern and SE rules.
@@
III-C Attack Point 3: Attack/Error in HMI Screen based on RTDB (Purple Dashed Line)
@@
III-C1 A Normal HMI Screen with Power Flow Analysis Using Only GenAI Without SoM Technique
The following “Prompt”-“Response” boxes show a normal HMI screen without any attack/error in segmented parts. Then, a prompt requested the GPT to arrange the segments based on its understanding. The process is that the 9 segments are given as inputs along with the following prompt, without any additional rules or principles to train the GPT model. As can be seen, the response demonstrates the incapability of the GPT model to arrange and make a connection between different segments of an HMI screen based on the 14-bus system in PowerWorld Simulator.
{promptbox}
Scenario #3A (GenAI without SoM)